OpenStudy (anonymous):

Which of the following sequences is not arithmetic? 1, 2, 3, 4, ... 2, 4, 6, 8, ... 0, -2, -4, -6, ... 1, 4, 9, 16, ...

OpenStudy (anonymous):

We can find out by checking each one, one at a time: (A) 2-1=1; 3-2=1, 4-3=1 Same difference each time, so (A) IS arithmetic. (B) 4-2=2; 6-4=2; 8-6=2 Same difference each time, so (B) IS arithmetic. (C) -2-0=-2; -4-(-2)=-2; -6-(-4)=-2 Same difference each time, so (C) IS arithmetic. (D) 4-1=3; 9-4=5; 16=9=7 NOT the same difference each time, so (D) IS NOT arithmetic.